OpenStudy (anonymous):

- 1/2= - 9/10 solve using the multiplication principle and simplify your answer.

OpenStudy (anonymous):

from where did you get this question?

OpenStudy (anonymous):

an assignment but I apologize I wrote it wrong. It should be - 1/2x = - 9/10

jhonyy9 (jhonyy9):

-10 = -18x --- x = -10/-18 = 5/9